# ASK: Adaptive Self-improving Knowledge Framework for Audio Text Retrieval

### v1

@@ 5.2 Main Results
We evaluate the effectiveness of our proposed ASK framework by integrating it into various baseline models and comparing their performance on the AudioCaps and Clotho datasets. The results are organized by the cross-modal interaction strategy.
@@
Table 1 presents the results for models employing a global, sentence-level interaction strategy. Our ASK framework demonstrates substantial and consistent improvements across both datasets and architectures. When applied to the ResNet-BERT baseline on AudioCaps, ASK improves the A2T R@1 score by a remarkable 6.0% absolute and the T2A R@1 score by 3.2% absolute . This strong performance gain validates the effectiveness of our core mechanisms in breaking the GLB and mitigating RDM. Furthermore, ASK proves to be model-agnostic, delivering significant gains on the more powerful transformer-based CED-SONAR architecture as well. For instance, on the challenging Clotho dataset, it boosts the A2T R@1 by up to 1.7% absolute and the T2A R@1 by 1.4% absolute. The results also highlight the flexibility of ASK in leveraging diverse knowledge sources, with different sources showing strengths on different dataset-architecture combinations.
@@
We also validate ASK on a strong baseline with a local, token-level interaction strategy (xie2024gpa). The Audio-to-Textretrieval results are presented in Table 2. The full results, including the Text-to-Audio retrieval scores, are detailed in Appendix D.
The results demonstrate that ASK delivers consistent and significant gains even on this fine-grained architecture. On AudioCaps, our best variant, ASK, improves the R@1 score by a substantial margin of 2.6% absolute. On the more challenging Clotho dataset, ASK+ achieves the top R@1 performance, boosting the baseline by 1.4% absolute. These improvements underscore the universal benefit of our framework; breaking the GLB and mitigating RDM are crucial enhancements regardless of whether the model’s interaction mechanism is global or local.
Table 2: Results for Audio-to-Text Retrieval under the local interaction strategy. The symbols +, , and denote different knowledge sources in Section 4.1.
AudioCaps Clotho
Method R@1 R@5 R@10 R@1 R@5 R@10
xie2024gpa $41.1_{\pm 0.3}$ $73.8_{\pm 0.4}$ $85.2_{\pm 0.6}$ $18.1_{\pm 0.2}$ $40.2_{\pm 0.3}$ $53.4_{\pm 0.4}$
@@
In addition to the in-domain evaluations, we conduct a challenging zero-shot cross-dataset experiment to further assess the generalization capabilities of ASK. The results, detailed in Appendix E, demonstrate that ASK significantly improves the model’s performance when transferring from AudioCaps to Clotho, confirming its strong generalization benefits.
5.3 Ablation Study and Analysis
@@ G. Method R@1 R@5 R@10 R@1 R@5 R@10
To validate the contribution of each component within our ASK framework, we conduct a series of ablation studies on the AudioCaps dataset using the ResNet-BERT architecture and an in-domain knowledge source (ASK+). The results are presented in Table 3.
@@
We first analyze the necessity of our multi-grained design. Removing the fine-grained knowledge base results in a substantial performance drop of 4.3% absolute in A2T R@1, confirming the critical role of instance-level details for precise retrieval. Similarly, removing the coarse-grained base leads to a 4.6% drop in A2T R@1, which underscores the importance of the global semantic prior provided by the prototypes. The full model, which leverages both, significantly outperforms either single-granularity variant, demonstrating that the fine- and coarse-grained knowledge sources are complementary.
@@
We then ablate the core mechanisms of ASK. 1) Knowledge Injection: Disabling the knowledge injection step causes a notable drop of 2.9% in A2T R@1. This empirically validates that creating gradient pathways to out-of-batch data is the primary driver for breaking the GLB and enhancing representations. 2) Reliability Weighting: Ablating our adaptive reliability weighting mechanism ("w/o Adaptive Reliability Weighting") results in a significant 2.7% drop in A2T R@1 and a 1.5% drop in T2A R@1. This provides strong evidence that not all retrieved knowledge is equally beneficial, and that modulating the loss based on cross-modal consistency is crucial for mitigating the impact of noises and achieving robust performance.
Figure 3: Ablation experiment on ASK+. Effect of the number $\mathcal{T}$ of Knowledge Update.
@@
To validate our solution to the Representation Drift Mismatch, we analyze the impact of the knowledge base update frequency $\mathcal{T}$. As shown in Table 3, removing the dynamic refinement entirely causes a substantial performance drop of 2.8% in A2T R@1. This provides strong empirical evidence for our theoretical analysis in Section 3.3, confirming that allowing the RDM to grow unchecked harms performance by injecting stale, misaligned knowledge.
Figure 3 further illustrates the sensitivity to $\mathcal{T}$. Performance initially improves as the update frequency increases, peaking at an optimal period of 15 epochs. This peak significantly outperforms both the static knowledge base and the baseline. However, updating too frequently also leads to suboptimal results. This suggests a trade-off: while frequent updates mitigate RDM, they may also disrupt the stability of the knowledge representation before the model has had sufficient time to learn from it. The results underscore the necessity of a co-evolving knowledge base and the importance of tuning the update frequency.
@@ Appendix A Derivation and Visualization of RDM’s Impact
We consider a simplified loss function $\mathcal{L}=\mathcal{L}_{\text{main}}(u_{i},u^{\prime}_{i})$ that incorporates a knowledge-enhanced representation $u^{\prime}_{i}=(1-\rho)u_{i}+\rho\mathcal{K}$, where $u_{i}=f_{\theta_{t}}(a_{i})$ and $\mathcal{K}$ is the expected representation of retrieved knowledge. The gradient of the loss with respect to the model parameters $\theta_{t}$ is:
@@
Bounding the Knowledge Deviation via RDM.
We now bound the norm of the deviation $\|\Delta\mathcal{K}\|_{2}$ using the RDM. We leverage Pinsker’s inequality (cover1999elements), which relates the KL divergence to the Total Variation Distance ($D_{TV}$):
@@ Conclusion.
Combining these steps, we have established a formal link: an increase in RDM (Eq. 4) widens the upper bound on the knowledge vector deviation $\|\Delta\mathcal{K}\|_{2}$ (Eq. 23), which in turn increases the potential magnitude of the gradient deviation $\Delta\nabla$ (Eq. 20). This increases the risk of gradient misalignment, which can lead to training instability. Our dynamic knowledge refinement mechanism is designed to mitigate this risk by periodically resetting the RDM to zero.
@@ Appendix B Theoretical Justification and Convergence of the ASK Objective
In this section, we provide a theoretical justification for the ASK framework. We demonstrate that our training procedure can be viewed as a principled alternating optimization algorithm designed to maximize the log-likelihood of the observed data, which in turn guarantees the monotonic non-increase of our final loss function and thus ensures convergence.
@@
We conceptualize our approach by introducing latent variables, $z_{i}=(z_{i,f},z_{i,c})$, representing the unobserved "optimal" knowledge for each sample $x_{i}$. The observed data likelihood is the marginal likelihood over these latent variables:
@@
$\mathcal{F}(Q,\theta_{i})$ $=\mathbb{E}_{Q(z_{i})}[\log p(x_{i},z_{i};\theta)]$ (28)
@@
  • We define the probability mass of these distributions over a specific neighbor $z_{j}$ using our reliability weights (Eq. 11):
@@
Our final modulated loss from Eq. 17,
@@
This two-stage alternating optimization guarantees that the total objective is non-decreasing at each full iteration, $\mathcal{L}(\theta_{t+1})\geq\mathcal{L}(\theta_{t})$. Consequently, minimizing the negative log-likelihood (our total loss $\mathcal{L}_{\text{ASK}}$) guarantees that the loss is monotonically non-increasing. Given that $\mathcal{L}_{\text{ASK}}$ is bounded below by zero, the Monotone Convergence Theorem ensures that the sequence of loss values converges to a limit, and the parameters $\{\theta_{t}\}$ converge to a stationary point
@@ Appendix C Optimal Transport for Batch-level Alignment
where $\langle\mathbf{Q},\mathbf{S}_{f}\rangle=\mathrm{tr}(\mathbf{Q}^{\top}\mathbf{S}_{f})$ is the total similarity score. $H(\mathbf{Q})=-\sum_{i,j}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ is the entropy regularizer, controlled by $\varepsilon>0$. The constraints enforce that the marginals of $\mathbf{Q}$ must sum to predefined distributions $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$, which represent the importance of each instance. Following prior work (su2017order), we set both $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ to a uniform distribution over the batch, i.e., $\frac{1}{|B|}\mathbf{1}_{|B|}$. This problem is efficiently solved for the optimal plan $\mathbf{Q}^{*}$ using the Sinkhorn-Knopp algorithm (cuturi2013sinkhorn).
@@ Appendix D Full Results for Local Interaction Strategy
This section provides the complete retrieval results for our experiments on the local, token-level interaction baseline, including both Audio-to-Text (A2T) and Text-to-Audio (T2A) directions. The main A2T results and their analysis are presented in the main paper. Table 4 presents the full comparison.
Table 4: Full results for Audio-Text Retrieval on AudioCaps and Clotho under the local interaction strategy. The symbols +, , and denote different knowledge sources in Section 4.1.
@@
As demonstrated in the second half of Table 4, the ASK framework consistently improves upon the baseline in the Text-to-Audio retrieval direction as well. On AudioCaps, ASK+ achieves the highest R@1 score, improving the baseline by 1.0% absolute. On Clotho, the ASK variant delivers the strongest R@1 performance with a significant gain of 1.2% absolute. These results confirm that the benefits of our proposed mechanisms are symmetric, enhancing both retrieval directions and validating the overall effectiveness of the ASK framework on fine-grained architectures.
Appendix E Zero-Shot Generalization
To further assess the generalization capabilities of our ASK framework, we conduct a zero-shot cross-dataset evaluation. In this setup, models are trained exclusively on the AudioCaps training set and then directly evaluated on the Clotho test set, without any fine-tuning. This challenging setting tests the model’s ability to generalize to a different data distribution. The results for the global ResNet-BERT architecture are presented in Table 5.
Table 5: Zero-shot generalization performance on the Clotho test set. All models were trained only on AudioCaps. The symbols +, , and denote different knowledge sources in Section 4.1.
A2T T2A
@@ Appendix E Zero-Shot Generalization
mei2022metric $12.8_{\pm 0.3}$ $29.0_{\pm 0.4}$ $39.7_{\pm 0.5}$ $10.1_{\pm 0.1}$ $27.6_{\pm 0.3}$ $38.3_{\pm 0.5}$
ASK $\textbf{14.1}_{\pm 0.2}$ $30.3_{\pm 0.4}$ $\textbf{43.7}_{\pm 0.5}$ $\textbf{11.9}_{\pm 0.1}$ $\textbf{31.2}_{\pm 0.4}$ $\textbf{42.8}_{\pm 0.4}$
ASK $13.6_{\pm 0.2}$ $30.2_{\pm 0.4}$ $40.5_{\pm 0.5}$ $\textbf{11.9}_{\pm 0.3}$ $30.5_{\pm 0.5}$ $40.3_{\pm 0.6}$
ASK+ $13.6_{\pm 0.2}$ $31.6_{\pm 0.4}$ $43.3_{\pm 0.5}$ $11.5_{\pm 0.3}$ $30.8_{\pm 0.5}$ $42.6_{\pm 0.5}$
The results clearly demonstrate that the ASK framework significantly enhances the model’s zero-shot generalization ability. All variants of ASK outperform the baseline across most metrics. Notably, the ASK variant, which leverages the large-scale, out-of-domain WavCaps dataset as its knowledge source, achieves the best overall performance. It improves the A2T R@1 by 1.3% absolute and the T2A R@1 by a substantial 1.8% absolute.
This finding is particularly insightful: by exposing the model to a diverse, external knowledge source during training, ASK equips the model with a richer, more robust semantic understanding. This allows it to better generalize to the unseen concepts and acoustic conditions present in the Clotho dataset, even when the primary training data is from AudioCaps. This validates that ASK is not merely an in-domain memorization technique, but a genuine framework for improving model generalization.

### v2

@@ 5.2 Main Results
We evaluate the effectiveness of our proposed ASK framework by integrating it into various baseline models. The results are organized by the cross-modal interaction strategy.
@@
Table 5 presents the results for models using a global, sentence-level interaction strategy. ASK consistently outperforms competitive methods across all datasets and architectures. On AudioCaps (ResNet-BERT), ASK surpasses the foundational ML-ACT by a remarkable 6.0% (A2T) and 3.2% (T2A) in absolute R@1. Crucially, it eclipses Auto-ACD, with further R@1 gains of up to 1.5% (ASK) and 2.2% (ASK+). These improvements validate our core mechanisms in successfully breaking the GLB and mitigating RDM. Demonstrating its model-agnosticism, ASK also significantly enhances the transformer-based CED-SONAR architecture. On the Clotho dataset, it outperforms the GLAP baseline, boosting A2T R@1 to 19.7% (+1.3%) and T2A R@1 to 16.3% (+1.2%). Finally, the varying optimal variants across setups highlight ASK’s flexibility in leveraging diverse knowledge sources.
Figure 3: Results for Audio-Text Retrieval on AudioCaps and Clotho under the local interaction strategy. The symbols +, , and denote different knowledge sources in Section 4.1.
@@
We also validate ASK with local, token-level interaction strategies (Xie et al., 2024; Wu et al., ). Figure 3 presents the retrieval results for our local, token-level interaction baselines. The results demonstrate consistent and significant gains across both retrieval directions, outperforming both the FLAM and GPA baselines. On AudioCaps, ASK achieves substantial improvements over the strongest baseline (GPA), with ASK enhancing the Audio-to-Text R@1 by 2.6% and ASK+ boosting the Text-to-Audio R@1 by 1.0%. On Clotho, ASK+ delivers the top Audio-to-Text R@1 performance, while ASK yields the best Text-to-Audio R@1. Notably, our variants exceed FLAM by even larger margins across all metrics. These symmetric improvements confirm the universal benefit of our framework.
@@
To strictly evaluate robustness, we conduct bidirectional zero-shot experiments: training on AudioCaps and testing on Clotho, and vice versa. As illustrated in Figure 4, the ASK framework consistently expands the retrieval performance envelope compared to the baseline across all metrics in both transfer directions. Specifically, when transferring from AudioCaps to Clotho (left), leveraging the diverse WavCaps knowledge source (ASK) yields a notable 1.3% absolute gain in A2T R@1. Similarly, in the Clotho-to-AudioCaps transfer (right), ASK variants maintain superior performance, with ASK achieving a significant improvement over the baseline (e.g., +2.6% in A2T R@1). These results confirm that our multi-grained knowledge injection effectively mitigates domain shifts and prevents overfitting to source-domain specifics.
Figure 4: Zero-shot performance on AudioCaps and Clotho. , , and + denote different knowledge sources in Section 4.1.
@@ G. Method R@1 R@5 R@10 R@1 R@5 R@10
5.3 Ablation Study and Analysis
To validate the contribution of each component within our ASK framework, we conduct a series of ablation studies on the AudioCaps dataset using the ResNet-BERT architecture and an in-domain knowledge source. The results are presented in Table 2.
@@ 5.3 Ablation Study and Analysis
We first analyze the necessity of our multi-grained design. Removing the fine-grained knowledge base results in a substantial performance drop of 4.3% absolute in A2T R@1, confirming the critical role of instance-level details for precise retrieval. Similarly, removing the coarse-grained base leads to a 4.6% drop in A2T R@1, which underscores the importance of the global semantic prior provided by the prototypes. The model, which leverages both, significantly outperforms either single-granularity variant, demonstrating that the fine- and coarse-grained knowledge sources are complementary.
Figure 5: Effect of the frequency $\mathcal{T}$ of Knowledge Update. Ablation experiment on ASK+ under the global interaction strategy.
@@
We then ablate the core mechanisms of ASK. 1) Knowledge Injection: Disabling the knowledge injection step causes a notable drop of 2.9% in A2T R@1. This empirically validates that creating gradient pathways to out-of-batch data is the primary driver for breaking the GLB and enhancing representations. 2) Reliability Weighting: Ablating our adaptive reliability weighting mechanism results in a significant 2.7% drop in A2T R@1 and a 1.5% drop in T2A R@1. This provides strong evidence that not all retrieved knowledge is equally beneficial, and that modulating the loss based on cross-modal consistency is crucial for mitigating the impact of noises and achieving robust performance.
@@
We evaluate the effect of the knowledge-base update period $\mathcal{T}$ on mitigating RDM. As shown in Table 2, disabling dynamic refinement leads to a 2.8% drop in A2T R1, empirically validating our theoretical claim in Section 3.3 that unchecked RDM introduces stale and misaligned knowledge.
Figure 5 shows that performance improves as the update frequency increases, reaching an optimum at $T=15$ epochs, which surpasses both the static knowledge base and the baseline. However, overly frequent updates degrade performance, indicating a trade-off: while frequent updates curtail RDM, they can also destabilize the knowledge representation before the model fully adapts. These findings highlight the necessity of a co-evolving knowledge base and careful tuning of the update frequency.
Table 3: Qualitative comparison of audio-text retrieval. GT represents Ground Truth.
Audio Query ML-ACT Mei et al. (2022) ASK (Ours)
[…]
GT: “Thunder roars in the distance as rain falls” (Confusing Texture) Top-1: “Food sizzling in a pan” (Audio Ambiguity) GT Rank: 21 Top-1: “Thunder roars in the distance as rain falls” GT Rank: 1
[…]
GT: “Church bells ringing” (Rare Event) Top-1: “Train horn blowing” (Tonal Similarity) GT Rank: 13 Top-1: “Church bells ringing” GT Rank: 1
Case Study.
Table 5.3 qualitatively illustrates how the ASK framework overcomes the Gradient Locality Bottleneck (GLB). In Scenario 1, the baseline model fails to resolve semantic ambiguity between acoustically similar textures like thunder and sizzling food, whereas ASK distinguishes these fine-grained details to achieve Rank 1. Likewise, for the long-tail Church bells concept, ASK avoids the baseline’s tendency to default to common tonal relatives like train horns. These results demonstrate that leveraging diverse out-of-batch knowledge during training allows ASK to learn a discriminative embedding space that generalizes to rare and ambiguous events without requiring explicit retrieval during inference.
@@ Appendix A Derivation and Visualization of RDM’s Impact
We consider a simplified loss function $\mathcal{L}=\mathcal{L}_{\text{main}}(u_{i},u^{\prime}_{i})$ that incorporates a knowledge-enhanced representation $u^{\prime}_{i}=(1-\rho)u_{i}+\rho\mathcal{K}$, where $u_{i}=f_{\theta_{t}}(a_{i})$ and $\mathcal{K}$ is the expected representation of retrieved knowledge. Formally, $\mathcal{K}=\sum_{j}P(j)z_{j}$, where $z_{j}$ denotes the embedding vector of the $j$-th sample in the knowledge base, and $P(j)$ is its corresponding probability weight. The gradient of the loss with respect to the model parameters $\theta_{t}$ is:
@@
Bounding the Knowledge Deviation via RDM
We now bound the norm of the deviation $\|\Delta\mathcal{K}\|_{2}$ using the RDM. We leverage Pinsker’s inequality, which relates the KL divergence to the Total Variation Distance ($D_{TV}$):
@@ Conclusion.
Combining these steps, we have established a formal link: an increase in RDM widens the upper bound on the knowledge vector deviation $\|\Delta\mathcal{K}\|_{2}$ (Eq. 26), which in turn increases the potential magnitude of the gradient deviation $\Delta\nabla$ (Eq. 23). This increases the risk of gradient misalignment, which can lead to training instability. Our dynamic knowledge refinement mechanism is designed to mitigate this risk by periodically resetting the RDM to zero.
@@ Appendix B Theoretical Justification and Convergence of the ASK Objective
In this section, we provide a complete theoretical justification for the ASK framework. We demonstrate that our training procedure can be viewed as a principled alternating optimization algorithm designed to maximize the log-likelihood of the observed data, which in turn guarantees the monotonic non-increase of our final loss function and thus ensures convergence.
@@
We conceptualize our approach by introducing latent variables, $z_{i}=(z_{i,f},z_{i,c})$, representing the unobserved optimal knowledge for each sample $x_{i}$. The observed data likelihood is the marginal likelihood over these latent variables:
@@
$\mathcal{F}(Q,\theta)$ $=\mathbb{E}_{Q(z_{i})}[\log p(x_{i},z_{i};\theta)]$ (31)
@@
  • We define the probability mass of these distributions over a specific neighbor $z_{j}$ using our reliability weights:
@@
Our final modulated loss,
@@
This two-stage alternating optimization guarantees that the total objective is non-decreasing at each full iteration, $\mathcal{L}(\theta_{t+1})\geq\mathcal{L}(\theta_{t})$. Consequently, minimizing the negative log-likelihood guarantees that the loss is monotonically non-increasing. Given that $\mathcal{L}_{\text{ASK}}$ is bounded below by zero, the Monotone Convergence Theorem ensures that the sequence of loss values converges to a limit, and the parameters $\{\theta_{t}\}$ converge to a stationary point
@@ Appendix C Optimal Transport for Batch-level Alignment
where $\langle\mathbf{Q},\mathbf{S}_{f}\rangle=\mathrm{tr}(\mathbf{Q}^{\top}\mathbf{S}_{f})$ is the total similarity score. $H(\mathbf{Q})=-\sum_{i,j}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ is the entropy regularizer, controlled by $\varepsilon>0$. The constraints enforce that the marginals of $\mathbf{Q}$ must sum to predefined distributions $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$, which represent the importance of each instance. Following prior work (Su and Hua, 2017), we set both $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ to a uniform distribution over the batch, i.e., $\frac{1}{|B|}\mathbf{1}_{|B|}$. This problem is efficiently solved for the optimal plan $\mathbf{Q}^{*}$ using the Sinkhorn-Knopp algorithm (Cuturi, 2013).
@@ Appendix D Full Results for Local Interaction Strategy
This section provides the complete retrieval results for our experiments on the local, token-level interaction baselines, including both Audio-to-Text and Text-to-Audio directions. Table 4 presents the full comparison against both the FLAM and GPA architectures.
Table 4: Full results for Audio-Text Retrieval on AudioCaps and Clotho under the local interaction strategy. The symbols +, , and denote different knowledge sources.
@@
As demonstrated in Table 4, ASK consistently outpaces both baselines across all metrics in the Text-to-Audio retrieval direction as well. When compared to the stronger GPA baseline, ASK+ achieves the highest R@1 score on AudioCaps, yielding a 1.0% absolute improvement, which translates to a substantial 2.0% margin over FLAM. On Clotho, the ASK variant delivers the strongest R@1 performance with a significant gain of 1.2% absolute over GPA, and a notable 2.5% absolute advantage over FLAM. These results confirm that the benefits of our proposed mechanisms are symmetric, enhancing both retrieval directions and validating the overall effectiveness of the ASK framework on fine-grained architectures.
Appendix E Comparison with Alternative Retrieval-Augmented Frameworks
To further validate the design choices of our ASK framework, we compare it against classical retrieval-augmented contrastive learning paradigms adapted from computer vision and information retrieval. Specifically, we benchmark against representative retrieval-augmented and contrastive learning paradigms: MoCo He et al. (2020), ANCE Xiong et al. , and NNCLR Dwibedi et al. (2021). These baselines represent the classical strategies for expanding negative sample capacity, performing hard negative mining, and utilizing external support sets, respectively. We adapt their core mechanisms to the Audio-Text Retrieval setting. All variants are built upon the identical ML-ACT Mei et al. (2022) backbone to ensure a fair and rigorous comparison.
Table 5: Comparison with other retrieval-augmented methods on AudioCaps. The symbol ++ denotes using training set as the knowledge source
Audio-to-Text Text-to-Audio
@@ Appendix E Comparison with Alternative Retrieval-Augmented Frameworks
ML-ACT Mei et al. (2022) $36.3_{\pm 0.5}$ $68.6_{\pm 0.3}$ $81.5_{\pm 0.2}$ $32.2_{\pm 0.4}$ $68.2_{\pm 0.1}$ $81.2_{\pm 0.2}$
NNCLR Dwibedi et al. (2021) $32.5_{\pm 0.9}$ $66.6_{\pm 0.5}$ $80.9_{\pm 0.7}$ $30.2_{\pm 0.3}$ $64.9_{\pm 0.1}$ $80.1_{\pm 0.1}$
ANCE Xiong et al. $35.9_{\pm 0.5}$ $67.8_{\pm 0.3}$ $83.6_{\pm 0.2}$ $31.7_{\pm 0.1}$ $65.4_{\pm 0.5}$ $79.8_{\pm 0.1}$
MoCo He et al. (2020) $37.2_{\pm 0.3}$ $68.5_{\pm 0.5}$ $82.0_{\pm 0.3}$ $32.9_{\pm 0.4}$ $67.4_{\pm 0.6}$ $80.9_{\pm 0.1}$
ASK+ $\textbf{42.0}_{\pm 0.2}$ $\textbf{74.2}_{\pm 0.5}$ $\textbf{85.4}_{\pm 0.6}$ $\textbf{35.4}_{\pm 0.3}$ $\textbf{70.2}_{\pm 0.3}$ $\textbf{83.1}_{\pm 0.7}$
As shown in Table 5, our proposed ASK significantly outperforms all alternative strategies across all metrics, achieving an R@1 of 42.0% in Audio-to-Text retrieval. Conversely, directly migrating classical contrastive methods to the Audio-Text Retrieval (ATR) domain yields sub-optimal or degraded performance. While MoCo provides only marginal improvements, both ANCE and NNCLR lead to noticeable performance drops, with NNCLR causing the most severe degradation (e.g., Audio-to-Text R@1 dropping from 36.3% to 32.5%).
We hypothesize that this performance degradation stems from the unique semantic ambiguities in ATR. Specifically, methods relying on strict hard negative mining (ANCE) or unimodal nearest neighbors (NNCLR) are highly susceptible to acoustic confusion, which inadvertently introduces false negatives or modality-specific noise into the contrastive objective. Additionally, the memory queues in MoCo may suffer from representation drift due to the continuous updating of the fine-grained audio encoder. Unlike these rigid mechanisms, ASK mitigates these domain-specific bottlenecks by employing an adaptive reliability weighting scheme that explicitly evaluates cross-modal consistency, effectively filtering out the retrieval noise that limits standard baselines.
Appendix F Hyperparameter Sensitivity Analysis
To comprehensively evaluate the robustness of the proposed ASK framework, we conduct sensitivity analyses on two critical hyperparameters: the retrieved knowledge size ($K$) and the knowledge injection ratio ($\rho$). The experiments are conducted on the AudioCaps dataset using the ASK+ variant. The results are illustrated in Figure 7.
Figure 7: Hyperparameter sensitivity analysis. Left: Impact of the retrieved knowledge size $K$. Right: Impact of the knowledge injection ratio $\rho$.
Impact of the Retrieved Knowledge Size ($K$).
The left panel of Figure 7 illustrates the effect of varying the number of retrieved knowledge samples $K$ from $5$ to $30$. The framework achieves its optimal performance in both Audio-to-Text (A2T) and Text-to-Audio (T2A) retrieval at $K=10$, reaching a peak A2T R@1 of 42.0% and T2A R@1 of 35.4%. When $K$ is set too small (e.g., $K=5$), the limited neighborhood fails to provide sufficient semantic diversity, restricting the model’s ability to generalize to ambiguous or long-tail concepts. Conversely, setting $K$ too large (e.g., $K\geq 15$) inevitably introduces acoustically similar but semantically irrelevant noise into the knowledge base. This excessive retrieval dilutes the valid semantic guidance and consequently degrades the alignment performance.
Impact of the Knowledge Injection Ratio ($\rho$).
The right panel of Figure 7 presents the performance fluctuations across different knowledge injection ratios $\rho\in[0.10,0.35]$. This hyperparameter controls the crucial trade-off between retaining the original instance identity and incorporating the retrieved global semantic prior. As shown in the figure, the model attains its peak performance at $\rho=0.20$. A lower injection ratio (e.g., $\rho=0.10$) provides insufficient global context, making it difficult to fully break the Gradient Locality Bottleneck (GLB). However, an excessively high ratio ($\rho\geq 0.25$) over-dominates the feature fusion process. This leads to feature homogenization, where the unique acoustic or textual characteristics of the original sample are overshadowed by the aggregated neighborhood representations, ultimately resulting in a decline in retrieval accuracy.
Appendix G Visualization of Adaptive Reliability Weighting
To explicitly address how the Adaptive Reliability Weighting mechanism mitigates Representation-Drift Mismatch (RDM) and filters noise, we delve into the mathematical behavior of the reliability potential $\mathcal{F}=-\log\Psi$. We provide an internal visualization of these assigned weights during the training process to demonstrate their dynamic modulation effects.
A critical but nuanced property of our framework is that $\mathcal{F}$ acts as a dynamic negative regularizer. Given the normalized embedding space, the cross-modal similarity generally maintains non-negative exponential characteristics, leading to an expected similarity $\Psi\geq 1$. Consequently, the reliability weight $\mathcal{F}$ consistently operates in the negative domain ($\mathcal{F}\leq 0$). This design intrinsically serves as a reward mechanism that dynamically relaxes the standard in-batch contrastive penalty when reliable external knowledge is injected.
Table 6 illustrates this internal mechanism using two challenging acoustic ambiguity scenarios. Query 1 involves confusing environmental textures (“Thunder” vs. “Sizzling food”). Query 2, drawn directly from our evaluation set, involves engine and wind noises where a “Motorboat” could be acoustically confused with a “Motorcycle”.
Table 6: Internal visualization of the reliability potential $\Psi$ and the negative regularizer weight $\mathcal{F}$.
Neighborhood State Semantic Concept $\Psi$ $\mathcal{F}$ Modulation
Query 1: “Thunder roars in the distance as rain falls” (Environmental Texture Ambiguity)
Clean Neighborhood “Loud thunder claps” 1.60 - 0.47 0.91
Noise Neighborhood “Food sizzling loudly” 1.05 - 0.05 0.99
Query 2: “A motorboat driving by as water splashes and wind blows” (Vehicle Engine Ambiguity)
Clean Neighborhood “A speedboat traveling across water” 1.55 - 0.44 0.91
Noise Neighborhood “A motorcycle engine revving” 1.08 - 0.08 0.98
When the retrieved neighborhood contains highly relevant concepts (e.g., “Loud thunder claps” or “A speedboat traveling”), the strong semantic alignment yields a larger $\Psi$ (>1>1), forcing $\mathcal{F}$ to be a distinctively negative value with a large magnitude (e.g., -0.47 and -0.44). This dynamically reduces the modulation multiplier $(1+\lambda\mathcal{F})<1$, effectively rewarding the model by down-weighting the standard contrastive penalty and encouraging the absorption of this clean knowledge.
Conversely, when the retrieved knowledge is corrupted by RDM drift—such as retrieving the acoustically similar but semantically incorrect “Food sizzling loudly” or “A motorcycle engine revving”—the weak cross-modal alignment results in a $\Psi$ closer to 1. This causes the negative regularizer $\mathcal{F}$ to approach zero (e.g., -0.05 and -0.08), which pulls the modulation multiplier back to near 1.0. By doing so, the framework withholds the reward and maintains the strict original contrastive loss, seamlessly preventing the model from over-relying on the drifted acoustic noise.